# RALLY: Role-Adaptive LLM-Driven Yoked Navigation for Agentic UAV Swarms

### v1

@@ II-B Problem Formulation
III Role-Adaptive LLM-driven Yoked Navigation
@@
III-A Two Stage LLM-Based Consensus Inference
The high-level consensus generation strategy $\text{LLM}_{\text{HC}}$ employs a two-stage structured prompting approach—local intention generation $\text{LLM}_{\text{init}}$ followed by neighborhood consensus refinement $\text{LLM}_{\text{cons}}$—to realize an LLM-driven personalized semantic decision mechanism. By integrating role definitions, threat analysis, and formation coverage requirements into natural language prompts, numerical observations are mapped into interpretable intentions. This design deeply couples the LLM’s semantic reasoning with various role logics. Example prompts are illustrated in Fig. 4.
@@ III-A Two Stage LLM-Based Consensus Inference
Firstly, each agent $i$ uses its current local observation $\bm{o}_{i}^{t}$ to generate initial intent through the LLM. Accordingly, we design a task-specific instruction $\text{X}_{\text{task}}$ that outlines agents’ mission requirements and game conditions, and craft the prompt ${\text{Y}_{\text{init}}}$ to capture the observation $\bm{o}_{i}^{t}$ of agent $i$. The LLM’s natural-language output is then parsed into numeric goals $\hat{g}_{i}^{t}$ as
$\hat{g}_{i}^{t}=\text{LLM}_{\text{init}}(\text{X}_{\text{task}},{\text{Y}_{\text{init}}}(\bm{o}_{i}^{t})).$ (5)
@@
$\bm{z}_{i}^{t}=\left(\bm{o}_{i}^{t},\left\{(\bm{p}_{j}^{t},\bm{v}_{j}^{t},\hat{g}_{j}^{t},k_{j}^{t})\right\}_{j\in\mathcal{N}_{i}^{t}}\right).$ (6)
On this basis, each agent $i$ constructs a new prompt ${\text{Y}_{\text{cons}}}$ describing its intents, $\text{RL}_{\text{HI}}$-produced role $k_{i}^{t}$, neighbors’ roles, as well as environmental constraints. We further incorporate task-driven Chain-of-Thought (CoT) prompts $\text{M}_{\text{CoT}}$, such as “Cluster or disperse based on the threats from enemy” and “needs cluster with other two teammates”, to strengthen the LLM’s reasoning and minimize hallucinations. The LLM is instructed to output a refined consensus like “I recommend going to target $[x,y]$” which is converted into the final numeric target:
@@
This refinement embeds a balance between maximizing swarm utility and ensuring safe enemy avoidance under adversarial and coverage requirements.
@@
For global state $\bm{s}^{t}$ and all agents’ goal decisions $\bm{g}^{t}:=(g_{1}^{t},\cdots,g_{N}^{t})$, the global value function $Q_{\text{tot}}$ is defined by a mixing network $f$, which will be detailed in III-B:
@@
Under Assmption 1, the two-stage policy achieves strictly higher expected return:
@@
For any $(\bm{o}_{i}^{t},k_{i}^{t})$, the two-stage consensus step selects:
$g_{i}^{*}=\underset{g\in\mathcal{T}}{\operatorname{argmax}}\ Q_{i}\left(\bm{o}_{i}^{t},k_{i}^{t},g_{i}^{t}\right),$ (12)
by construction of $\text{LLM}_{\text{cons}}$. In contrast, the one-stage policy outputs $\tilde{g}_{i}=\widetilde{\text{LLM}}_{\text{HC}}(\bm{o}_{i}^{t},k_{i}^{t})$ without refinement. By optimality of $g_{i}^{*}$:
$Q_{i}(\bm{o}_{i}^{t},k_{i}^{t},g_{i}^{*})\geq Q_{i}(\bm{o}_{i}^{t},k_{i}^{t},\tilde{g}_{i}),$ (13)
with strict inequality when consensus improves the goal. Under Assmption 1, increasing any individual agent’ value $Q_{i}$ cannot decrease the total value, which implies:
$Q_{\text{tot}}(\bm{s}^{t},\bm{g}^{*})\geq Q_{\text{tot}}(\bm{s}^{t},\tilde{\bm{g}}),$ (14)
@@
Theorem 1 shows that under Assumption 1, the two-stage policy $\bm{\pi}^{\text{two}}_{H}=(\text{LLM}_{\text{cons}}\circ\text{LLM}_{\text{init}})$ achieves strictly higher expected cumulative reward than the one-stage baseline $\bm{\pi}^{\text{one}}_{H}=\widetilde{\text{LLM}}_{\text{HC}}$. Next, we introduce role-based credit-assignment mechanism to ensure that Assumption 1 always holds.
III-B Credit-Assignment Mechanism Based on Role-value Mixing Network
@@
22:           Compute target values as follows: $y^{j}&=R^{j}+\gamma_{\text{rmix}}\;\max_{\bm{k}^{\prime}}\bar{Q}\bigl{(}\bm{s}^{j+1},\bm{k}^{\prime};\bar{\theta}\bigr{)};\end{aligned}$
@@
$k_{i}^{t}=\underset{k_{i}^{t}\in\mathcal{K}}{\operatorname{argmax}}\ Q_{i}\left(\bm{o}_{i}^{t},k_{i}^{t}|\hat{g}_{i}^{t}\right).$ (15)
By Eq. (8), in dynamic multi-agent adversarial tasks, instantaneous reward $R$ cannot directly reflect the contribution from the choice of role $k_{i}^{t}$. To address this issue, we propose an RMIX-assisted credit-assignment mechanism to evaluate each agent’s role choices. Notably, considering the significant inference latency of LLM, which results in higher costs for acquiring training samples, we introduce a semi-offline learning approach to improve sample efficiency. Similar to Curriculum Learning (CL)[52], we first exploit the zero-shot capability of LLM (e.g., GPT-4o[53]) via API to obtain its role assignment strategy $\bm{\pi}_{\text{GPT}}$. By interacting with the environment under local observation $\bm{o}_{i}^{t}$, the mid-layer policy $\bm{\pi}_{M}$ and the physical flight controller $\bm{\pi}_{L}$, we record GPT’s preliminary role suggestions ${k}_{i_{\text{GPT}}}$ and store the related transition quadruple $\langle{\bm{o}}_{i},{k}_{i_{\text{GPT}}},\bm{s}^{\prime},R\rangle$ into a replay buffer $\mathcal{B}_{R}$, where $\bm{s}^{\prime}$ denotes the transitioned state. This offline data collection runs in parallel with LLM consensus reasoning, enabling RMIX to assist the LLM in understanding role assignments and reducing the cold-start overhead from ineffective exploration.
@@
Subsequently, we proceed to the standard CTDE online training. RMIX jointly learns from both offline and newly collected online samples to continuously explore and refine cooperative behaviors among agents. Generally, RMIX can be solved by a classical Stochastic Gradient Descent (SGD) approach with the cost function of RMIX being formulated as:
$L(\theta)=\sum_{i=1}^{|\mathcal{B}_{R}|}\left[\left(y_{i}^{\text{tot}}-Q_{\text{tot}}(\bm{s}^{t},k^{t};\theta,\hat{g}^{t})\right)^{2}\right],$ (17)
@@
$y^{\text{tot}}=R^{t}+\gamma_{\text{rmix}}\max_{k^{\prime}}\bar{Q}\left(\bm{s}^{\prime},k^{\prime};\bar{\theta},\hat{g}^{t}\right),$ (18)
@@
III-C Context-Based LLM Fine-Tuning
@@
7:  $\mathcal{B}_{\text{fil}}\!\leftarrow\!\{x\in\mathcal{B}_{\text{LLM}}\mid\text{Check}(x)\geq 1\}$ cf. Eq. (20);
@@
Albeit the proficiency of LLM like GPT-4o[53], a lightweight version is preferred by resource-constrained multi-agent architecture. Therefore, through self-generated instruction tuning [54], we introduce a capacity-migration augmentation, so as to improve a smaller model’s reasoning ability. Through a concerted effort, we successfully transfers task-understanding capabilities from a State-Of-The-Art (SOTA) foundation model to a local version, and subsequently compresses the model to under 5GB of memory usage, thus potentially enabling distributed inference of the consensus reasoning module on onboard UAV GPUs[55].
@@
where $g_{i}^{*}$ denotes the manually labeled decision, and $\mathscr{I}_{i}^{*}$ constitutes additional consensus inference outputs, stored beyond Eq. (7), specifically for training purposes. In particular, to guide the generation of desired output in Eq. (7), human-annotated instructions, similar to those in Fig. 4 and 10, are provided. On this basis, GPT-4o’s output ${g}_{i}$ serves as RALLY’s consensus decision for selecting the agent’s target region. We then interact with the mid-layer policy $\bm{\pi}_{M}$ and the physical flight controller $\bm{\pi}_{L}$ to obtain the next local observation $\bm{o}_{i}^{t+1}$ and role $k_{i}^{t+1}$. This process repeats until acquiring a dataset $\mathcal{B}_{\text{LLM}}$ with sufficient inference samples and trajectory data $(\bm{o}_{i},k_{i},g_{i}^{*},\mathcal{I}_{i}^{*},R)$
@@
where $\mathbb{I}\left(\cdot\right)$ is an indicator function that returns $1$ when the condition is established and 00 otherwise; $w_{1,g}$, $w_{2,g}$, $w_{3,g}$, $w_{4,g}$ are the indicator importance weights; $\Lambda$ is a pre-defined set of anomalous characters; $L_{\min}$ and $L_{\max}$ denote the minimum and maximum inference result lengths, respectively.
@@
where $\psi$ denotes the low-rank adaptation parameters. To reduce computational and storage requirements, we use the LLaMA-Factory framework[56] with LoRA [57] to fine-tune a smaller, $\psi$-parameterized LLM (e.g., Qwen2.5 or Llama3.1).
@@
IV Experimental Results and Discussions
IV-A Experimental Settings
TABLE II: Key Environmental and Reward Function Parameter Configurations
@@ IV-A Experimental Settings
UAV velocity range (m/s) [$-1$, $1$]
Adversary velocity range (m/s) [$-0.75$ , $0.75$]
Observation distance $\delta_{\text{obs}}$ (m) $3$
@@
In this section, we evaluate the effectiveness and superiority of RALLY on DS-CEFC tasks in MPE[47] and a fully distributed, high-fidelity SITL[48] simulation built on Gazebo–ROS–PX4. Notably, each simulation runs for $1,000$ steps, wherein every $50$ steps, UAVs select their target regions according to $\pi_{H}$. The set of possible targets is $\mathcal{G}_{a}=\{(p_{x},p_{y})|p_{x},p_{y}\in\{-8,0,8\}\}$ with both coordinates initialized uniformly in $[-8,8]$ m and re-sampled when a region’s urgency reaches zero. The urgency level $\kappa_{\text{tr}}^{t}$ is initialized as $1$. Whenever there exist a formation of UAV covering the region $\text{tr}\in\mathcal{T}$, $\kappa_{\text{tr}}^{t}$ will decrease linearly with a scale $\omega_{d}$ until zero; otherwise, it remains unchanged. Other key environment and mission parameters are listed in Table II.
Figure 6: Performance comparison between RALLY and other methods.
@@
We compare RALLY against three baselines, CIHRL[2], CoNavGPT[28], and DITTO[49], to evaluate its effectiveness on the DS-CEFC task. CIHRL [2] incorporates multi-agent communication and belongs to the SOTA MARL approach for DS-CEFC. CoNavGPT [28] employs an LLM as a global planner without any training process, achieving astonishing success rates and efficiency on the navigation task. DITTO [49] achieves good collaboration based on LLM, demonstrating strong role-based heterogeneity. All non-API LLMs are deployed in a distributed fashion on $8$ NVIDIA GeForce RTX 4090 GPUs (24 GB each).
@@
IV-B Performance Comparison with Baselines
IV-B1 Overall Performance

### v2

@@ 2.2 Problem Formulation
3 Role-Adaptive LLM-driven Yoked Navigation
@@
3.1 Two Stage LLM-Based Consensus Inference
The high-level consensus generation strategy $\text{LLM}_{\text{HC}}$ employs a two-stage structured prompting approach—local intention generation $\text{LLM}_{\text{init}}$ followed by neighborhood consensus refinement $\text{LLM}_{\text{cons}}$—to realize an LLM-driven personalized semantic decision mechanism. By integrating role definitions, threat analysis, and formation coverage requirements into natural language prompts, numerical observations are mapped into interpretable intentions. This design deeply couples the LLM’s semantic reasoning with various role logics. Notably, the LLM reasoning currently does not rely on historical memory and solely on the given context.
@@ 3.1 Two Stage LLM-Based Consensus Inference
Firstly, each agent $i$ uses its current local observation $\bm{o}_{i}^{t}$ to generate initial intent through the LLM. Accordingly, we design a task-specific instruction $\text{X}_{\text{task}}$ that outlines agents’ mission requirements and task conditions, and craft the prompt $\text{Y}_{\text{init}}$} to capture the observation $\bm{o}_{i}^{t}$ of agent $i$. The LLM’s natural-language output is then parsed into numeric goals ${g^{\prime}}_{i}^{t}$ as
${g’}_{i}^{t}=\text{LLM}_{\text{init}}(\text{X}_{\text{task}},{\text{Y}_{\text{init}}}(\bm{o}_{i}^{t}),\text{M}_{\text{CoT}}).$ (5)
Next, the Role Selection Policy $\text{RL}_{\text{HI}}$ takes the current observation $\bm{o}_{i}^{t}$ as input to optimize the role selection for each agent. The final selected role ${k}_{i}^{t}$ is given as
${k}_{i}^{t}=\text{RL}_{\text{HI}}(\bm{o}_{i}^{t};{g^{\prime}}_{i}^{t}),$ (6)
which determines the agent’s action strategy based on its role.
@@
$\bm{z}_{i}^{t}=\left(\bm{o}_{i}^{t},\left\{(\bm{p}_{j}^{t},\bm{v}_{j}^{t},{g^{\prime}}_{j}^{t},k_{j}^{t})\right\}_{j\in\mathcal{N}_{i}^{t}}\right).$ (7)
On this basis, each agent $i$ constructs a new prompt $\text{Y}_{\text{cons}}$} describing its intents, $\text{RL}_{\text{HI}}$-produced role $k_{i}^{t}$, neighbors’ roles, as well as environmental constraints. We further incorporate task-driven Chain-of-Thought (CoT) prompts $\text{M}_{\text{CoT}}$, such as “Cluster or disperse based on the threats from enemy” and “needs cluster with other two teammates”, to strengthen the LLM’s reasoning and minimize hallucinations. The LLM is instructed to output a refined consensus like “I recommend going to target [x,y][x,y]” which is converted into the final numeric target:
@@
This refinement embeds a balance between maximizing swarm utility and ensuring safe avoidance of enemy under adversarial and coverage requirements.
Example prompts for $\text{X}_{\text{task}}$ and $\text{M}_{\text{CoT}}$ are illustrated in Fig. 4, while more prompt details, such as $\text{Y}_{\text{init}}$ and $\text{Y}_{\text{cons}}$, are provided in the Appendix. Notably, as evidenced in Fig. 19 of the Appendix, the design of prompts shows considerable sensitivity to the environmental changes and yields significantly different responses.
@@
For global state $\bm{s}^{t}$ and all agents’ goal decisions $\bm{g}^{t}:=(g_{1}^{t},\cdots,g_{N}^{t})$, the global value function $Q_{\text{tot}}$ is defined by a mixing network $f$, which will be detailed in Section III.3.2:
@@
Assumption 2 (Performance Improvement of Extra Contextual Reasoning).
Similar to [54, 55, 56], the extra contextual reasoning helps guide the decision-making process toward more robust and effective outcomes, which leads to higher-quality $Q$-values.
@@
Under Assumption 1 and Assumption 2, the two-stage policy achieves strictly higher expected return:
@@
Under Assumption 2, for any $(\bm{o}_{i}^{t},k_{i}^{t})$, the two-stage consensus step selects
$Q_{i}(\bm{o}_{i}^{t},k_{i}^{t},g_{i}^{t})\geq Q_{i}(\bm{o}_{i}^{t},k_{i}^{t},{g^{\prime}}_{i}^{t}),\forall{g^{\prime}}_{i}^{t}$ (13)
by construction of $\text{LLM}_{\text{cons}}$. In contrast, the one-stage policy outputs $\tilde{g}_{i}=\widetilde{\text{LLM}}_{\text{HC}}(\bm{o}_{i}^{t},k_{i}^{t})$ without refinement. In other words, Assumption 2 implies
$Q_{i}(\bm{o}_{i}^{t},k_{i}^{t},g_{i}^{t})\geq Q_{i}(\bm{o}_{i}^{t},k_{i}^{t},\tilde{g}_{i}^{t}).$ (14)
Under Assumption 1, increasing any individual agent’ value $Q_{i}$ cannot decrease the total value, which implies:
[…]
@@
Theorem 1 shows that under Assumptions 1 and 2, the two-stage policy $\bm{\pi}^{\text{two}}_{H}=(\text{LLM}_{\text{cons}}\circ\text{LLM}_{\text{init}})$ achieves strictly higher expected cumulative reward than the one-stage baseline $\bm{\pi}^{\text{one}}_{H}=\widetilde{\text{LLM}}_{\text{HC}}$. In other words, while it may not guarantee the optimal goal decision, the two-stage consensus process provides a more comprehensive evaluation than the initial decision, potentially leading to an improved target decision. Next, we introduce role-based credit-assignment mechanism to ensure that Assumption 1 always holds.
3.2 Credit-Assignment Mechanism Based on Role-value Mixing Network
@@
22:    Compute target values as follows: $\begin{aligned} y^{j}&=R^{j}+\gamma_{\text{rmix}}\;\max_{\bm{k}^{\prime}}\bar{Q}\bigl{(}\bm{s}^{j+1},\bm{k}^{\prime};\bar{\theta}\bigr{)};\end{aligned}$
@@
$k_{i}^{t}=\underset{k_{i}^{t}\in\mathcal{K}}{\operatorname{argmax}}\ Q_{i}\left(\bm{o}_{i}^{t},k_{i}^{t};{g^{\prime}}_{i}^{t}\right).$ (16)
By Eq. (9), in dynamic multi-agent adversarial tasks, instantaneous reward $R$ cannot directly reflect the contribution from the choice of role $k_{i}^{t}$. To address this issue, we propose an RMIX-assisted credit-assignment mechanism to evaluate each agent’s role choices. Also, we consider the significant inference latency of LLM, which results in higher costs for acquiring training samples. In particular, with $8$ agents and $3$ roles each, the dimension of joint role space turns $3^{8}$, and when combined with sparse, adversarial rewards, this makes pure online exploration extremely difficult. Therefore, we resort to an offline learning approach to improve sample efficiency. Similar to Curriculum Learning (CL)[57], we first exploit the zero-shot capability of LLM (e.g., GPT-4o[58]) via API to obtain its role assignment strategy $\bm{\pi}_{\text{GPT}}$. To clarify, $\bm{\pi}_{\text{GPT}}$ refers to the online GPT4o-API version, rather than any locally fine-tuned large model. By interacting with the environment under local observation $\bm{o}_{i}^{t}$, the mid-layer policy $\bm{\pi}_{M}$ and the physical flight controller $\bm{\pi}_{L}$, we record GPT’s preliminary role suggestions ${k}_{i_{\text{GPT}}}$ and store the related transition quadruple $\langle{\bm{o}}_{i},{k}_{i_{\text{GPT}}},\bm{s}^{\prime},R\rangle$ into a replay buffer $\mathcal{B}_{R}$, where $\bm{s}^{\prime}$ denotes the transitioned state. This offline data collection runs in parallel with LLM consensus reasoning and seeds the replay buffer with sensible role assignments, enabling RMIX to assist the LLM in understanding role assignments and reducing the cold-start overhead from ineffective exploration.
@@
Subsequently, we proceed to the standard CTDE online training. RMIX jointly learns from both offline and newly collected online samples to continuously explore and refine cooperative behaviors among agents. Generally, RMIX can be solved by a classical Stochastic Gradient Descent (SGD) approach, with the cost function of RMIX being formulated as:
$L(\theta)=\sum_{i=1}^{|\mathcal{B}_{R}|}\left[\left(y_{i}^{\text{tot}}-Q_{\text{tot}}(\bm{s}^{t},k^{t};\theta,{g}^{t})\right)^{2}\right],$ (18)
@@
$y^{\text{tot}}=R^{t}+\gamma_{\text{rmix}}\max_{k^{\prime}}\bar{Q}\left(\bm{s}^{\prime},k^{\prime};\bar{\theta},{g}^{t}\right),$ (19)
@@
3.3 Context-Based LLM Fine-Tuning
@@
[…]
@@
Albeit the proficiency of LLMs like GPT-4o[58], a lightweight version is preferred by resource-constrained multi-agent system. Therefore, through self-generated instruction tuning [59], we introduce a capacity-migration augmentation, so as to improve a smaller model’s reasoning ability. Through a concerted effort, we successfully transfer task-understanding capabilities from a State-Of-The-Art (SOTA) foundation model to a local version, and subsequently compress the model to under 5GB of memory usage, thus potentially enabling distributed inference of the consensus reasoning module on onboard UAV GPUs[60].
@@
where $g_{i}^{*}$ denotes the manually labeled decision, and $\mathscr{I}_{i}^{*}$ constitutes additional consensus inference outputs, stored beyond Eq. (7), specifically for training purposes. In particular, to guide the generation of desired output in Eq. (8), human-annotated instructions, similar to those in Fig. 4 and 11, are provided. On this basis, GPT-4o’s output ${g}_{i}$ serves as RALLY’s consensus decision for selecting the agent’s target region. We then interact with the mid-layer policy $\bm{\pi}_{M}$ and the physical flight controller $\bm{\pi}_{L}$ to obtain the next local observation $\bm{o}_{i}^{t+1}$ and role $k_{i}^{t+1}$. This process repeats until acquiring a dataset $\mathcal{B}_{\text{LLM}}$ with sufficient inference samples and trajectory data $(\bm{o}_{i},k_{i},g_{i}^{*},\mathcal{I}_{i}^{*},R)$.
@@
where $\mathbb{I}\left(\cdot\right)$ is an indicator function that returns $1$ when the condition is established and 0 otherwise; $w_{1,g}$, $w_{2,g}$, $w_{3,g}$, $w_{4,g}$ are the indicator importance weights; $\Lambda$ is a pre-defined set of anomalous characters; $L_{\min}$ and $L_{\max}$ denote the minimum and maximum inference result lengths, respectively.
@@
where $\psi$ denotes the low-rank adaptation parameters. To reduce computation and storage requirements, we use the LLaMA-Factory framework[61] with LoRA [62] to fine-tune a smaller, $\psi$-parameterized LLM (e.g., Qwen2.5 or Llama3.1).
@@
4 Experimental Results and Discussions
4.1 Experimental Settings
Table 2: Key Environmental and Reward Function Parameter Configurations
@@ 4.1 Experimental Settings
UAV velocity range (m/s) [1-1, 11]
Adversary velocity range (m/s) [0.75-0.75 , $0.75$]
Observation distance $\delta_{\text{obs}}$ (m) 33
@@
Figure 6: Performance comparison between RALLY and baselines.
In this section, we evaluate the effectiveness and superiority of RALLY on DS-CEFC tasks in MPE[49] and a fully distributed, high-fidelity SITL[50] simulation built on Gazebo–ROS–PX4. Concretely, each simulation runs for $1,000$ steps, wherein every $50$ steps, UAVs select their target regions according to $\pi_{H}$. The set of possible targets is $\mathcal{G}_{a}=\{(p_{x},p_{y})|p_{x},p_{y}\in\{-8,0,8\}\}$ with both coordinates initialized uniformly in [8,8][-8,8] m and re-sampled when a region’s urgency reaches zero. The urgency level $\kappa_{\text{tr}}^{t}$ is initialized as $1$. Whenever there exists a formation of UAVs covering the region $\text{tr}\in\mathcal{T}$, $\kappa_{\text{tr}}^{t}$ will decrease linearly with a scale $\omega_{d}$ until zero; otherwise, it remains unchanged. The pursuer’s speed is set to be twice that of the evader to ensure it has the necessary mobility to chase down and capture the evaders effectively. Other key environment and mission parameters are listed in Table 2.
@@
We compare RALLY against three representative baselines, CIHRL[2], CoNavGPT[30], and DITTO[51], to evaluate its effectiveness on the DS-CEFC task. CIHRL [2], which does not incorporate role assignment, incorporates multi-agent communication and belongs to the SOTA MARL approach for DS-CEFC. CoNavGPT [30] employs an LLM as a global planner without any training process, achieving high success rates and efficiency on the navigation task. DITTO [51] achieves good collaboration based on LLM, demonstrating strong role-based heterogeneity. All the non-API LLMs are deployed in a distributed fashion on $8$ NVIDIA GeForce RTX 4090 GPUs (24 GB each).
Figure 10: Generalization of RALLY to varying target areas.
@@
4.2 Performance Comparison with Baselines
4.2.1 Overall Performance